### v1

# OoDIS: Anomaly Instance Segmentation Benchmark

Abstract
Autonomous vehicles require a precise understanding of their environment to navigate safely. Reliable identification of unknown objects, especially those that are absent during training, such as wild animals, is critical due to their potential to cause serious accidents. Significant progress in semantic segmentation of anomalies has been driven by the availability of out-of-distribution (OOD) benchmarks. However, a comprehensive understanding of scene dynamics requires the segmentation of individual objects, and thus the segmentation of instances is essential. Development in this area has been lagging, largely due to the lack of dedicated benchmarks. To address this gap, we have extended the most commonly used anomaly segmentation benchmarks to include the instance segmentation task. Our evaluation of anomaly instance segmentation methods shows that this challenge remains an unsolved problem. The benchmark website and the competition page can be found at: https://vision.rwth-aachen.de/oodis.
[…]
1 Introduction
Figure 1: Annotation example for the previous semantic annotation of the RoadAnomaly21 dataset (top) and the extended annotation labels (bottom) for our newly proposed benchmark.
Modern segmentation methods [8, 7] perform well on curated closed-world datasets with a fixed set of classes. However, models trained with a fixed training set fall short of solving the task when unexpected objects are present [17, 18]. These anomalies often cause models to misclassify, assigning known classes to unknown objects [15, 21]. To prevent such behavior in real world applications, it is important to design or adapt models to handle such anomalies. The task of anomaly detection spans multiple modalities [36, 3, 30, 27], applications [2, 24], and tasks [11, 37, 35]. The particular focus of this work is the anomaly instance segmentation task, that aims to provide segmentation models with the ability to segment out-of-distribution (OOD) objects. This task is particularly critical for autonomous driving scenarios, where a recognition error can cause serious accidents. A collision with lost cargo on the road or with livestock could be life-threatening. To evaluate the performance of anomaly segmentation methods, a number of benchmarks have been proposed [31, 5].
While anomaly segmentation [28, 35, 25] methods achieve exciting results on popular benchmarks, the area of anomaly instance segmentation remains unexplored. Early datasets [31] for anomaly segmentation included partial instance annotations of anomalies, but recently proposed datasets omit instance information [4, 5]. However, instance segmentation is critical for understanding complex scenes with multiple anomalous objects, such as cows and sheep as shown in Figure 1, that may appear in a group. Previous anomaly segmentation approaches that operate on a pixel level would fail to distinguish individual objects. Understanding these objects separately provides context about the potential dynamics of a scene, improving downstream tasks such as navigation or planning. We hypothesize that recent advances in open set [36, 20] and class-agnostic [22] instance segmentation have encouraged research in the area of anomaly instance segmentation, which was previously too challenging. Recently, three works following different paradigms proposed to solve the task of anomaly instance segmentation [12, 32, 29]. However, each of these works proposes a different evaluation procedure.
To address this limitation, we propose a benchmark and evaluate existing methods in a unified manner. We extend the labels of popular anomaly segmentation datasets [4, 5] to instance segmentation. These datasets provide diverse real-world cases of road anomalies with precise annotations. We reuse the Average Precision (AP) metric [16] for instance evaluation similarly to the Cityscapes setup [9], with a slight modification to evaluate instances as small as 10 pixels in size. In comparison to the semantic anomaly benchmarks, the AP metric avoids size bias and requires high precision for smaller anomalous objects. This is particularly important in the context of autonomous driving, where detecting anomalies in the distance is critical to give the system time to react.
To this end, we re-annotated anomalies within the Fishyscapes [4], RoadAnomaly21, and RoadObstacle21 [5] datasets to evaluate anomaly instance segmentation methods. We apply publicly available instance segmentation methods on both validation and test set and provide qualitative evaluation of the results. Our evaluations show that while current anomaly segmentation methods perform well on semantic anomaly segmentation, instance segmentation methods achieve moderate performance, suggesting a considerable space for improvement. We make validation data available on our challenge website, and open a submission portal where new approaches can be submitted.
2 Related Work
Out-of-Distribution (OOD) Datasets  have primarily focused on classification tasks, with several benchmarks recently introduced [37, 39]. A common evaluation task is disentanglement of two classification datasets such as CIFAR and SVHN. Methods such as deep ensembles [23] and Monte Carlo dropout [34], while performing well on OOD classification, show limited usefulness in anomaly segmentation tasks [5]. Open-set instance segmentation [20, 36] assumes the presence of OOD data during training, a condition not applicable to anomaly segmentation where completely unseen objects may appear [12]. In autonomous driving, novel evaluation schemes have been proposed for detection tasks [11, 24]. However, these works do not address the need for precise pixel-level mapping in monocular driving detection setups. Our work explores the segmentation of anomaly instances, which allows accurate prediction of individual, previously unseen, objects.
Anomaly Segmentation Datasets. Anomaly segmentation has received significant attention with the emergence of several recent datasets and benchmarks [31, 4, 5]. The Lost and Found (L&F) dataset [31] introduced the task of anomaly segmentation in a camera setup similar to the one used for the Cityscapes dataset [9]. L&F has annotations limited to the road area and anomaly classes; however, it has questionable labels that include bicycles and kids as anomalies [4]. To fully control for anomalies in the training and test sets, the CAOS benchmark [19] introduces a real dataset based on BDD100K [38], treating certain inlier classes as anomalies, and a synthetic dataset for training and testing. FishyScapes Lost and Found (FS L&F) [4] reannotates images from L&F to extend in-distribution regions outside of the road class and introduces a separate benchmark with artificial anomalies. Despite its popularity, FS L&F lacks anomaly instance segmentation and it is constrained to lost cargo on the road. To solve the diversity issue, SegmentMeIfYouCan [5] introduces a diverse dataset with real anomalies on roads, which are not limited to the Cityscapes camera perspective. In past years, evaluation on FS L&F and SegmentMeIfYouCan dataset has been a standard practice. However, instance annotations are missing from these datasets. Our work aims to extend these popular benchmarks by providing accurate instance annotations.
Anomaly Segmentation Methods. Segmentation of anomaly instances has been underexplored until recently. There are previous works in open-set instance segmentation [36, 20]. However, they rely on unknown objects present in the training set; and methods that rely on depth cues [33] that are not applicable in general case. In general anomaly instance segmentation methods produce per-pixel anomaly scores, while providing anomaly instances too. U3HS [12] uses uncertainty in semantic predictions to guide the region segmentation, and then clusters predicted class-agnostic instance embeddings. Mask2Anomaly [32] applies modifications to the Mask2Former [8] architecture to produce reliable semantic anomaly scores in background regions, and uses a connected components on anomaly scores with a strategy to remove false-positives using intersections with in-distribution predictions. UGainS [29] combines the RbA anomaly segmentation method [28] with an interactive segmentation model [22] to predict instances using point prompting. Given the limited number of specialized methods for anomaly instance segmentation, we evaluate these models and analyze their performance, offering insights into their practical applications and limitations.
Table 1: Evaluation of three existing anomaly segmentation methods. We observe improved performance when using extra networks and extra out-of-distribution (OOD) data. However, low scores suggests significant potential for improvement on our benchmark.
Method OOD Data Extra Network FishyScapes RoadAnomaly21 RoadObstacle21 Mean
AP AP50 AP AP50 AP AP50 AP AP50
UGainS [29] 27.14 45.82 11.42 19.15 27.22 46.54 25.19 42.81
Mask2Anomaly [32] 11.73 23.64 04.78 09.03 17.23 28.44 13.73 24.30
U3HS [12] 00.19 00.73 00.00 00.00 00.22 00.62 00.19 00.58
3 Benchmark Design.
Anomaly segmentation as a task attempts to identify unexpected objects unknown during training. Common examples include a deer or a cardboard box that may appear in the middle of the road. Per-pixel segmentation does not provide sufficient information for downstream tasks such as tracking or navigation. The more challenging problem of instance segmentation remains under-explored and lacks accessible benchmarks. This benchmark addresses the lack of test evaluation protocols available to the community.
We aim to fill the gap by extending the labels of SegmentMeIfYouCan [5] and FS L&F [4] datasets for instance segmentation. We merge these datasets into a unified benchmark and adopt commonly used Average Precision (AP) metrics [26], that closely follows the Cityscapes [9] segmentation benchmark.
Data. We use three datasets for anomaly segmentation: RoadAnomaly21 and RoadObstacle21 from SegmentMeIfYouCan [5], and FS L&F [4]. These are the standard benchmarks for the task, and they complement each other in label diversity well (see Figure 2). To maintain data integrity, we keep the test sets from the datasets intact, using $100$ images from RoadAnomaly21, $412$ from RoadObstacle21, and $275$ from FS L&F as our full test sets. In addition, we provide a relabeled validation set of 100 images from FS L&F.
The test set contains three relabeled datasets with different properties, but shares a common in-distribution dataset. For the submission to the benchmark, we allow models trained on $19$ Cityscapes [9] classes as the in-distribution dataset, and allow the use of auxiliary data, such as COCO [26] to introduce virtual anomalies, similar to other anomaly segmentation works [13, 28, 35, 10, 6, 14]. It is important to note that we expect no explicit supervision to segment unknowns, much like in the real world, we do not know what kind of anomalies we will encounter. The benchmark data contains three classes: inlier, outlier, and ignore. In-distribution regions contain classes known to Cityscapes; ignore regions are ambiguous regions that neither contain anomalies nor are in-distribution regions; and the outlier class contains anomalous instances (see Figure 1). Ignore regions are ambiguous regions for which a class cannot be defined; common cases in Cityscapes are: bridges, advertisement posts, back side of street signs and dark regions where the class could not be determined. We omit ignore regions in evaluation and discard cases that overlap significantly with these regions. We evaluate predictions only for the outlier class, without focusing on evaluation of in-distribution predictions. To calculate the final Average Precision (AP) score, we compute a weighted average based on the number of images in each dataset.
Labeling Policy. In RoadAnomaly21, anomalies are of arbitrary size, located anywhere on the image, containing highly diverse samples. Each individual object, such as an animal or object, is labeled as an individual object without introducing group labels. FS L&F mainly contains anomalies on the road, separate objects such as stacked boxes, which are treated as separate instances. Only ambiguous regions are treated as ignore for RoadAnomaly21 and FS L&F. For RoadObstacle21, however, only the drivable area is considered an inlier, and everything outside the drivable area, including anomalies, are labeled as ignore regions. Gaps within complex anomalies are also treated as ignore regions. Each labeled object on an image is given a unique identifier. Bounding boxes are also generated to facilitate anomaly localization.
Metrics. Conventional anomaly segmentation metrics tend to favor larger objects. Average Precision or False Positive Rate (FPR) per-pixel metrics, or sIoU, which groups anomalies together, do not provide the correct evaluation metric. Our benchmark uses the Average Precision (AP) metric, a standard in instance segmentation that evaluates precision at IoU thresholds from $0.5$ to $0.95$. Additionally, we provide the AP50 metric to assess performance at a 50% IoU threshold, following the community practice.
Detection Benchmark. While our current focus is instance segmentation, we have converted instance data and predictions into bounding boxes to evaluate anomaly object detection capabilities. However, our initial results show that current anomaly detection methods such as VOS [11] perform suboptimally in this setup. For more details on the detection benchmark we refer readers to the supplementary material and leave this area for future research.
4 Evaluated Methods & Discussion of Results
We evaluate existing anomaly instance segmentation methods (see Table 1). To ensure correctness, we contacted authors of the original works, and asked them for a submission to the benchmark. In cases when code was not available, we worked closely with authors to reimplement unavailable methods and submit them to the benchmark. We kept the test set private and allowed evaluation on the validation set.
The U3HS [12] method belongs to a class of models that neither require auxiliary data nor external models for instance segmentation. The core of the method is the ability to learn class-agnostic instance embeddings that generalize beyond the training distribution. These embeddings in uncertain regions are clustered to get instance predictions. This allows clustering of anomalous regions occluded by other objects. While U3HS is capable of localizing anomaly instances without external data, it struggles in generating precise object masks, as measured by the AP metric that evaluates instances with at least $50\%$ IoU with the ground truth.
[…]
Figure 2: Diversity of instance labels. RodAnomaly21  typically contains multiple objects, while RoadObstacle21  contains smaller objects in smaller quantities, and Fishyscapes L&F  provides a balance between the two.
Mask2Anomaly [32] is a model that uses auxilary data, but does not use an external model for instance segmentation. Common to other methods in the community [13, 35], the model uses auxiliary data from COCO [26] for guiding the anomaly scores that are grouped using connected components to form instance proposals. To reduce the number of false positives, Mask2Anomaly introduces a post-processing strategy. It computes the intersection with predicted in-distribution masks and uses class entropy to determine true instance proposals. The approach benefits from a powerful backbone and is effective in segmenting individual anomalous objects, however, it merges closely located anomalies (see Figure 3).
UGainS [29] is a method that uses both auxiliary data and an external generalist segmentation model, namely the segment anything model (SAM) [22]. The method uses the anomaly segmentation method RbA [28] based on Mask2Former [8], fine-tuned using data from COCO, to generate uncertainty regions. UGainS uses farthest point sampling to sample a number of points from these regions as prompts for SAM [22]. While the method produces accurate segmentation masks, it relies on two models to get predictions. A limited number of prompts leads to missed detections in smaller regions and increases the number of false positives in other areas. However, it demonstrates strong performance and produces well-separated instance masks.
(a) Ground Truth
(b) UGainS
(c) Mask2Anomaly
(d) U3HS
Figure 3: Qualitative comparison of the methods. The scene contains multiple grouped anomaly objects close to the camera and multiple smaller instances in the distance.
5 Conclusion
Detecting and accurately segmenting anomaly instances on roads is a significant challenge, requiring an understanding of ’objectness’ without direct training on specific anomaly classes. In this work, we introduced a new benchmark for anomaly instance segmentation that integrates three popular anomaly datasets. The unified benchmark provides a diverse set of anomalies that vary in size, number of images, and annotation detail. We evaluate the performance of current methods for segmenting anomaly instances and provide intuition behind the results. Our results show that current techniques struggle particularly with distant and small objects, and with precise segmentation masks. The benchmark results suggest strong opportunities for advancement in the area. As autonomous vehicle technologies continue to evolve, driven by large amounts of data, it remains a challenge to capture all possible real-world situations. Our work addresses the need to evaluate instance segmentation as a step towards reliable autonomous driving.
Acknowledgment
We acknowledge fruitful discussions with H. Blum, R. Chan, S. Gasperini and S. Rai, as well as a contribution of annotations for Fishyscapes by H. Blum, and help with the benchmark submission from S. Gasperini and S. Rai. M.R. an M.A. acknowledge support by the German Federal Ministry of Education and Research within the junior research group project “UnrEAL” (grant no. 01IS22069). Alexey Nekrasov’s research has been funded by BMBF project ”WestAI” (grant no. 01IS22094D). We thank M. Burdorf, G. Lydakis, C. Schmidt, and others in the lab for discussions and feedback.
Supplementary Material
Detection benchmark.
We have converted instance labels into bounding boxes for the anomaly detection benchmark. For evaluation, we considered three methods, namely UGainS [29], Mask2Anomaly [32], and VOS [11]. The COCO [26] Average Precision (AP) and Average Recall (AR) metrics serve as evaluation metrics. Unfortunately, we observed an unexpectedly poor performance of VOS. While performing well on ambiguous objects, i.e. the toy car is correctly predicted as an anomaly, vos struggles to predict for an unknown object (see Figure 4). Note that, we have not contacted the authors of VOS for help with the submission and cannot fully trust our results. We plan to open the detection benchmark for submission along with the instance benchmark, such that we can evaluate anomaly detection methods with the help of the community.
Qualitative Results.
We provide additional qualitative results in Figure 5.
Competition and Benchmark Website.
We follow a setup common [1] for hosting the benchmark. We host competition webpage (see Figure 6) on https://codalab.lisn.upsaclay.fr/ servers, and a benchmark webpage on our local server, with manually updated leaderboard for methods with at least an arXiv paper (see Figure 7).
Figure 4: VOS prediction on the Lost and Found dataset.
(a) Label
(b) U3HS
(c) Mask2Anomaly
(d) UGainS
Figure 5: Qualitative results on FS L&F, RoadAnomaly21 and RoadObstacle21 dataset.
Figure 6: Competition website overview.
Figure 7: Leaderboard on the website.