### v2

# OoDIS: Anomaly Instance Segmentation and Detection Benchmark

Abstract
Safe navigation of self-driving cars and robots requires a precise understanding of their environment. Training data for perception systems cannot cover the wide variety of objects that may appear during deployment. Thus, reliable identification of unknown objects, such as wild animals and untypical obstacles, is critical due to their potential to cause serious accidents. Significant progress in semantic segmentation of anomalies has been facilitated by the availability of out-of-distribution (OOD) benchmarks. However, a comprehensive understanding of scene dynamics requires the segmentation of individual objects, and thus the segmentation of instances is essential. Development in this area has been lagging, largely due to the lack of dedicated benchmarks. The situation is similar in object detection. While there is interest in detecting and potentially tracking every anomalous object, the availability of dedicated benchmarks is clearly limited. To address this gap, this work extends some commonly used anomaly segmentation benchmarks to include the instance segmentation and object detection tasks. Our evaluation of anomaly instance segmentation and object detection methods shows that both of these challenges remain unsolved problems. We provide a competition and benchmark website under https://vision.rwth-aachen.de/oodis.
[…]
I Introduction
Modern segmentation methods [1, 2] as well as object detection methods [3] perform well on curated closed-world datasets with a fixed set of classes. However, models trained with a fixed training set fall short of solving the task when unexpected objects are present [4, 5]. These anomalies often cause models to misclassify, assigning known classes to unknown objects [6, 7]. To prevent such behavior in real world applications, it is important to design or adapt models to handle such anomalies. The task of anomaly detection spans multiple modalities [8, 9, 10, 11], applications [12, 13], and tasks [14, 15, 16]. The particular focus of this work is two-fold. We focus on 1) the anomaly instance segmentation task, that aims to provide segmentation models with the ability to segment out-of-distribution (OOD) objects, and closely related to that 2) the anomaly object detection task, that aims to provide bounding boxes for OOD objects. Both tasks are particularly critical for autonomous driving scenarios, where a recognition error can cause serious accidents. A collision with lost cargo on the road or with livestock could be life-threatening. To evaluate the performance of anomaly segmentation methods, a number of benchmarks have been proposed [17, 18].
Figure 1: Annotation example for the previous semantic annotation of the RoadAnomaly21 dataset (top) and the extended annotation labels (bottom) for our newly proposed benchmark. Accurate instance level bounding boxes and segmentation masks are provided as part of OoDIS to evaluate methods for anomaly instance segmentation as well as anomalous object detection.
While anomaly segmentation [19, 16, 20] methods achieve exciting results on popular benchmarks, the area of anomaly instance segmentation remains unexplored. Early datasets [17] for anomaly segmentation included partial instance annotations of anomalies, but recently proposed datasets omit instance information [21, 18]. However, instance segmentation is critical for understanding complex scenes with multiple anomalous objects, such as cows and sheep as shown in Figure 1, that may appear in a group. Previous anomaly segmentation approaches that operate on a pixel level would fail to distinguish individual objects. Understanding these objects separately provides context about the potential dynamics of a scene, improving downstream tasks such as navigation or planning. We hypothesize that recent advances in open set [8, 22] and class-agnostic [23] instance segmentation have encouraged research in the area of anomaly instance segmentation, which was previously too challenging. Recently, three works following different paradigms proposed to solve the task of anomaly instance segmentation [24, 25, 26]. However, each of these works proposes a different evaluation procedure. The situation is similar in object detection. An in-distribution (ID) and OOD combination of datasets used for the evaluation of anomalous object detection is PascalVOC and [27] and COCO [28]. The classes available in COCO but not in PascalVOC serve as OOD classes, see [14]. In anomaly object detection, a challenging and unified evaluation procedure will be beneficial for the development of the field.
To address the outlined limitations, we propose two benchmarks and evaluate existing anomalous instance segmentation and object detection methods in a unified manner, respectively. We extend the labels of popular anomaly segmentation datasets [21, 18] to instance segmentation, which naturally includes bounding boxes. These datasets provide diverse real-world cases of road anomalies with precise annotations. We reuse the Average Precision (AP) metric [29] for instance evaluation similarly to the Cityscapes setup [30], with a slight modification to evaluate instances as small as 10 pixels in size. In comparison to the semantic anomaly benchmarks, the AP metric in both tasks avoids size bias and requires high precision for smaller anomalous objects. This is particularly important in the context of autonomous driving, where detecting anomalies in the distance is critical to give the system time to react.
To this end, we re-annotated anomalies within the Fishyscapes [21], RoadAnomaly21, and RoadObstacle21 [18] datasets to evaluate anomaly instance segmentation and anomalous object detection methods. We apply publicly available instance segmentation and object detection methods on both validation and test set and provide quantitative evaluation of the results. Our evaluations show that while current anomaly segmentation methods perform well on semantic anomaly segmentation, instance segmentation methods only achieve moderate performance, suggesting a considerable space for improvement. In anomalous object detection, currently available methods exhibit a clear deficiency in detecting anomalous objects in challenging environmental conditions. This holds also true for more recent open-vocabulary object detectors such as [31]. We make validation data available on our challenge website, and open a submission portal where new approaches can be submitted.
II Related Work
Out-of-Distribution (OOD) Datasets  have primarily focused on classification tasks, with several benchmarks recently introduced [15, 32]. A common evaluation task is disentanglement of two classification datasets such as CIFAR and SVHN. Methods such as deep ensembles [33] and Monte Carlo dropout [34], while performing well on OOD classification, show limited usefulness in anomaly segmentation tasks [18]. Open-set instance segmentation [22, 8] assumes the presence of OOD data during training, a condition not applicable to anomaly segmentation where completely unseen objects may appear [24]. In autonomous driving, novel evaluation schemes have been proposed for detection tasks [14, 13]. However, these works do not address the need for precise pixel-level mapping in monocular driving detection setups. Our work explores the segmentation of anomaly instances, which allows accurate prediction of individual, previously unseen, objects.
Anomaly Segmentation Datasets. Anomaly segmentation has received significant attention with the emergence of several recent datasets and benchmarks [17, 21, 18]. The Lost and Found (L&F) dataset [17] introduced the task of anomaly segmentation in a camera setup similar to the one used for the Cityscapes dataset [30]. L&F has annotations limited to the road area and anomaly classes; however, it has questionable labels that include bicycles and kids as anomalies [21]. To fully control for anomalies in the training and test sets, the CAOS benchmark [35] introduces a real dataset based on BDD100K [36], treating certain inlier classes as anomalies, and a synthetic dataset for training and testing. FishyScapes Lost and Found (FS L&F) [21] reannotates images from L&F to extend in-distribution regions outside of the road class and introduces a separate benchmark with artificial anomalies. Despite its popularity, FS L&F lacks anomaly instance segmentation and it is constrained to lost cargo on the road. To solve the diversity issue, SegmentMeIfYouCan [18] introduces a diverse dataset with real anomalies on roads, which are not limited to the Cityscapes camera perspective. In past years, evaluation on FS L&F and SegmentMeIfYouCan dataset has been a standard practice. However, instance annotations are missing from these datasets. Our work aims to extend these popular benchmarks by providing accurate instance annotations.
Anomaly Segmentation Methods. Segmentation of anomaly instances has been underexplored until recently. There are previous works in open-set instance segmentation [8, 22]. However, they rely on unknown objects present in the training set; and methods that rely on depth cues [37] that are not applicable in general case. In general anomaly instance segmentation methods produce per-pixel anomaly scores, while providing anomaly instances too. U3HS [24] uses uncertainty in semantic predictions to guide the region segmentation, and then clusters predicted class-agnostic instance embeddings. Mask2Anomaly [25] applies modifications to the Mask2Former [1] architecture to produce reliable semantic anomaly scores in background regions, and uses a connected components on anomaly scores with a strategy to remove false-positives using intersections with in-distribution predictions. UGainS [26] combines the RbA anomaly segmentation method [19] with an interactive segmentation model [23] to predict instances using point prompting. Given the limited number of specialized methods for anomaly instance segmentation, we evaluate these models and analyze their performance, offering insights into their practical applications and limitations.
Anomalous Object Detection Datasets & Methods. Up to now, compared to the amount of existing anomaly segmentation benchmarks, the availability of dedicated ID-OOD dataset combinations for object detection is limited. The CODA benchmark [13] introduces corner case datasets that contain rare scenes that naturally occur during driving. With about 90% of the recorded corner cases claimed as novel. Compared to the CODA benchmark, our benchmark proposed in the present work contains rather remote OOD objects, atypical for street scenes, while CODA contains rather near OOD objects that can be found comparatively frequently.
In [14] (VOS), PascalVOC [27] is used as ID and complemented with COCO [28], where the classes from COCO that are not in PascalVOC serve as OOD classes. Everyday scenes are not safety-critical per-se. In [38] (SAFE), BDD100k serve as ID and COCO as well as PascalVOC as OOD, however the task in [38] is to find OOD data rather than OOD objects. In light of these experimental setups, the lack of dedicated ID-OOD dataset combinations in safety-critical contexts becomes apparent, which is addressed by the present work. In a recent work [39] based on a preprint of the present work [40] it turned out that open-vocabulary object detectors such as grounding Dino [31] show clearly superior performance over conventional anomalous object detection methods such as VOS. However, it also turns out that there is still plenty of room for further improvement. A summary of the related dataset landscape is provided in Tab. I.
TABLE I: Comparison to previous anomaly benchmarks. The number of connected components are stated with respect to the test set. (*): For the CODA dataset, about 90% of the instances are assessed to be anomalous in [13].
Dataset Year Size Anomaly No. of Sequences Instance Bounding Private Benchmark
(Test/Val) Source Components (#) Labels Boxes
Wuppertal OoD Tracking
Street Obstacle Sequences 2022 1129 / - Staged Obj. 1592 ✓(20) instance tracking
Carla Wildlife 2022 1210 / - Simulated / Staged Obj. 2826 ✓(26) instance tracking
Wuppertal Obstacle Sequences 2022 938 / - Staged Obj. 1039 ✓(44) instance tracking
CODA Corner Case Dataset
CODA-KITTI 2022 309 / - Natural Occurance 399(∗) subsampled object detection
CODA-nuScenes 2022 134 / - Natural Occurance 1125(∗) subsampled object detection
CODA-ONCE 2022 1057 / - Natural Occurance 4413(∗) subsampled object detection
Fishyscapes
FS Lost & Found 2019 275 / 100 Staged Obj. 165(?) subsampled semantic segmentation
SegmentMeIfYouCan
RoadAnomaly21 2021 100 / 10 Web Sourcing 262 semantic segmentation
RoadObstacle21 2021 412 / 30 Staged Obj. 388 subsampled semantic segmentation
OoDIS
OoDIS 2024 787 / 140 Combined 1735 subsampled instance segmentation
& ✗ & object detection
III Benchmark and Evaluated Methods
Anomaly segmentation as a task attempts to identify unexpected objects unknown during training. Common examples include a deer or a cardboard box that may appear in the middle of the road. Per-pixel segmentation does not provide sufficient information for downstream tasks such as tracking or navigation. The challenging problems of instance segmentation and object detection remain under-explored and lack accessible benchmarks. This benchmark addresses the lack of test evaluation protocols available to the community.
We aim to fill the gap by extending the labels of SegmentMeIfYouCan [18] and FS L&F [21] datasets for instance segmentation, which naturally includes object detection. We merge these datasets into a unified benchmark and adopt commonly used Average Precision (AP) metrics [28], that closely follows the Cityscapes [30] segmentation benchmark.
Data. We use three datasets for anomaly segmentation and detection: RoadAnomaly21 and RoadObstacle21 from SegmentMeIfYouCan [18], and FS L&F [21]. These are the standard benchmarks for the task, and they complement each other in label diversity well (see Figure 2). To maintain data integrity, we keep the test sets from the datasets intact, using $100$ images from RoadAnomaly21, $412$ from RoadObstacle21, and $275$ from FS L&F as our full test sets. In addition, we provide a relabeled validation set of 100 images from FS L&F.
The test set contains three relabeled datasets with different properties, but shares a common in-distribution dataset. For the submission to the benchmark, we allow models trained on $19$ Cityscapes [30] classes as the in-distribution dataset, and allow the use of auxiliary data, such as COCO [28] to introduce virtual anomalies, similar to other anomaly segmentation works [41, 19, 16, 42, 43, 44]. It is important to note that we expect no explicit supervision to segment unknowns, much like in the real world, it is a priori unknown know what kind of anomalies a given vision system will encounter.
The benchmark data contains three classes: inlier, outlier, and ignore. In-distribution regions contain classes known to Cityscapes; ignore regions are ambiguous regions that neither contain anomalies nor are in-distribution regions; and the outlier class contains anomalous instances (see Figure 1). Ignore regions are ambiguous regions for which a class cannot be defined; common cases in Cityscapes are: bridges, advertisement posts, back sides of street signs and dark regions where the class could not be determined. We omit ignore regions in evaluation and discard cases that overlap significantly with these regions. We evaluate predictions only for the outlier class, without focusing on evaluation of in-distribution predictions. To calculate the final Average Precision (AP) score, we compute a weighted average based on the number of images in each dataset.
Altogether, our instance segmentation versions of Fishyscapes L&F, RoadAnomaly21 and RoadObstacle21 provide a diverse setting for numerical experiments. The objects vary considerably in their size and, in particular in RoadAnomaly, the scenes have strongly varying numbers of instances, see Fig. 2.
[…]
Figure 2: Diversity of instance labels. RodAnomaly21  typically contains multiple objects, while RoadObstacle21  contains smaller objects in smaller quantities, and Fishyscapes L&F  provides a balance between the two.
Labeling Policy. In RoadAnomaly21, anomalies are of arbitrary size, located anywhere on the image, containing highly diverse samples. Each individual object, such as an animal or object, is labeled as an individual object without introducing group labels. FS L&F mainly contains anomalies on the road, separate objects such as stacked boxes, which are treated as separate instances. Only ambiguous regions are treated as ignore for RoadAnomaly21 and FS L&F. For RoadObstacle21, however, only the drivable area is considered an inlier, and everything outside the drivable area, including anomalies, are labeled as ignore regions. Gaps within complex anomalies are also treated as ignore regions. Each labeled object on an image is given a unique identifier. Bounding boxes are also generated to facilitate anomaly localization and allow for the evaluation of anomaly object detection.
Metrics. Conventional anomaly segmentation metrics tend to favor larger objects. Per pixel Average Precision (AP) or False Positive Rate (FPR) metrics, or sIoU [18], which groups anomalies together, do not provide the correct evaluation metric. We utilize the Cityscapes instance segmentation [30] evaluation suite for anomaly instance segmentation as well as the COCO evaluation suite [28] for anomalous object detection. Our benchmark primarily uses the AP metric (not the be confused with the per-pixel AP), a standard in instance segmentation and object detection. We describe it briefly in the following:
[…]
As additional evaluation metric, we provide the average recall (AR). Restricting the number of predictions to $k$, only keeping predictions with highest confidence $s$, the corresponding recall $\mathrm{REC}t(k)$ is averaged over a chosen set $T$ of IoU thresholds, i.e., $\mathrm{AR}k=\frac{1}{|T|}\sum_{t\in T}\mathrm{REC}t(k)$. We provide AR1, AR10 and AR100. The recall reflects how many of the anomalous objects can be segmented / detected and how many of them are overlooked. Finally, we also report the number of the Predictions Per Frame (PPF) averaged over all images of the respective dataset, inspired by [46]. This metric detects an over-production of predictions, e.g. aiming at matching the ground truth by chance.
Methods for Anomaly Instance Segmentation and Anomalous Object Detection.
The U3HS [24] method belongs to a class of models that neither require auxiliary data nor external models for instance segmentation. The core of the method is the ability to learn class-agnostic instance embeddings that generalize beyond the training distribution. These embeddings in uncertain regions are clustered to get instance predictions. This allows clustering of anomalous regions occluded by other objects.
Mask2Anomaly [25] is a model that uses auxilary data, but does not use an external model for instance segmentation. Common to other methods in the community [41, 16], the model uses auxiliary data from COCO [28] for guiding the anomaly scores that are grouped using connected components to form instance proposals. To reduce the number of false positives, Mask2Anomaly introduces a post-processing strategy. It computes the intersection with predicted in-distribution masks and uses class entropy to determine true instance proposals.
UGainS [26] is a method that uses both auxiliary data and an external generalist segmentation model, namely the segment anything model (SAM) [23]. The method uses the anomaly segmentation method RbA [19] based on Mask2Former [1], fine-tuned using data from COCO, to generate uncertainty regions. UGainS uses farthest point sampling to sample a number of points from these regions as prompts for SAM [23].
TABLE II: Evaluation of three existing anomaly segmentation methods for three different datasets in terms of AP and AP50. Higher scores indicate stronger performance.
Method OOD Data Extra Network FishyScapes RoadAnomaly21 RoadObstacle21 Mean
AP AP50 AP AP50 AP AP50 AP AP50
UGainS [26] 27.14 45.82 11.42 19.15 27.22 46.54 25.19 42.81
Mask2Anomaly [25] 11.73 23.64 04.78 09.03 17.23 28.44 13.73 24.30
U3HS [24] 00.19 00.73 00.00 00.00 00.22 00.62 00.19 00.58
TABLE III: Evaluation of the three methods from Table II w.r.t. anomalous object detection. The detection results are split over three datasets.
Method OOD Data Extra Network FishyScapes RoadAnomaly21 RoadObstacle21 Mean
AP AP50 AP AP50 AP AP50 AP AP50
UGainS [26] 15.55 24.27 11.69 17.48 08.05 11.54 11.14 16.75
Mask2Anomaly [25] 01.01 02.33 01.43 02.91 01.34 01.99 01.24 02.23
U3HS [24] 00.31 00.75 00.04 00.17 00.09 00.22 00.16 00.40
IV Experiments
In this section we present quantitative results obtained by U3HS, Mask2Anomaly and UGainS on OoDIS’ new annotations provided for the FS Lost & Found, RoadAnomaly21 and RoadObstacle21. The results for all three methods were computed using the original code repositories provided by the respectively publications, if available. To ensure correctness, we contacted authors of the original works, and asked them for a submission to the benchmark. We received a submission for Mask2Anomaly by the authors. In case of U3HS the code was not available. We worked closely with the authors to reimplement it and submit the results to the benchmark. During this process, we kept the test sets of OoDIS private. The authors used the validation set for debugging and parameter tuning. In all experiments presented, we stick to default hyperparameters of the methods. We used low score thresholds to obtain rather high numbers of predictions. The PPF counts provide a measure for this.
In this section, we first present a comparative study of the three methods on the three datasets for both tasks, anomaly instance segmentation and anomalous object detection, focusing on the popular metrics AP and AP50. Thereafter we proceed with a more detailed analysis, providing results for further evaluation metrics. This is complemented with a choice of qualitative results.
Anomaly Instance Segmentation. In Tab. II we provide anomaly instance segmentation results for all three methods and all three datasets in terms of AP and AP50. The three methods vary greatly in performance. In terms of AP, none of them perform strongly, demonstrating that there is a room for further method development. The AP50 scores for all of the data subsets are below 47%, indicating that accurate localization on all three datasets is challenging. A closer comparison of results on the different datasets reveals that RoadAnomaly21 is particularly challenging. This can be explained as follows. In RoadObstacle21, the number of connected components in the annotations increased from 388 for the semantic segmentation version to 557 instances for instance segmentation. Most of the scenes contain one to two objects. However, in RoadAnomaly21 this count considerably increased from 262 to 739. The number of objects per scenes varies more strongly, with extreme cases where a flock of sheep consists of several dozens of instances. These cases are particularly hard to segment accurately. The segment sizes a summarized in Fig. 2.
Anomalous Object Detection. In Tab. III, an evaluation analogously to the previous paragraph on anomaly instance segmentation is provided for three methods over three datasets in terms of AP and AP50. Compared to the results for anomalous instance segmentation, two key differences can be observed. On the one hand, while the ranking of the methods is the same, a pronounced reduction in the AP and AP50 scores is observed. Smaller deviations of an instance segmentation, e.g. at the border of a component, are not punished by the segmentation IoU as much as for the object detection IoU. In the latter metric, only slight deviations can have a strong effect. On the other hand, while for anomalous instance segmentation the results on RoadAnomaly21 were considerably worse than on the other datasets, this is not the case anymore for anomalous object detection. Here, the lowest scores are obtained on RoadObstacle21. The reason for this might be due to the objects in RoadObstacle21 being smaller than in RoadAnomaly21. Accurate localization becomes increasingly challenging when considering smaller objects. In summary, the task of accurately localizing anomalous objects can be deemed a challenging task. Note that for practical purposed such as automated driving or robotic control, localization accuracy matters.
TABLE IV: Evaluation of three existing anomaly segmentation methods on the detection benchmark. The numbers are computed as weighted averages over all three datasets, where the weighting takes the datasets’ sizes into account.
Method OOD Data Extra Network
[…]
UGainS [26] 11.14 16.75 14.98 39.07 41.45 12.36
Mask2Anomaly [25] 01.24 02.23 00.09 18.87 19.45 09.74
U3HS [24] 00.16 00.40 01.08 01.82 01.83 03.80
LabelU3HSMask2AnomalyUGainS Fishyscapes L&F Road Anomaly21 Road Obstacle21 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 3: Qualitative comparison of all three methods on all three datasets.
Additional Evaluation Metrics. In Tab. IV, we provide additional evaluation metrics for anomalous object detection. All performance metrics are averaged over the three datasets. The AR1, A10 and AR100 scores additionally indicate that it is particularly challenging for all the tested methods to detect the anomalous objects present in the data. The majority of objects remain overlooked, also for the most strongly performing methods, namely UGainS. Noteworthily, the small difference between AR10 and AR100 shows that all methods can hardly deal with the very crowded scenes of the dataset such as flocks of sheep on the roads.
TABLE V: Evaluation of three existing anomaly instance segmentation methods for three different buckets of object sizes in terms of AP and AP50. Higher scores indicate stronger performance.
Object Size UGainS [26] M2A [25] U3HS [24]
AP AP50 AP AP50 AP AP50
$<1,\!000$ pix 8.50 19.94 1.32 2.68 0.15 0.43
$1,\!000$-$10,\!000$ pix 34.43 55.32 30.29 53.24 0.08 0.31
$>10,\!000$ pix 17.41 31.69 32.19 48.25 0.05 0.15
Mean 25.19 42.81 13.73 24.30 0.19 0.58
Evaluation for Different Object Sizes. In Tab. V we present instance segmentation results of the three methods, averaged over all three datasets for three different size ranges. This evaluation reveals that small objects are particularly difficult to find for all three methods. From a method perspective, it is noteworthy that UGainS and Mask2Anomaly are on par in medium sized object from 1.000–10.000 pixels, and Mask2Anomaly is even slightly superior for large objects. However, it gets clearly outperformed on the small objects below 1.000 pixels, which constitute the most challenging cases of the benchmark. Note that all mean values in the benchmark are computed as weighted averages where the weighting treats all images across all datasets uniformly. Hence, the size-related results do not average out to the provided mean results.
Qualitative Results. Figure 3 provides a qualitative comparison of the three methods on all three datasets. Across all three datasets it can be observed that UH3S is able to spot the anomalous objects, however it is unable to provide accurate segmentation masks, which leads to the low numbers on the benchmarks. Mask2Anomaly already provides strong segmentation performance, however, it can be seen from the qualitative examples that some instances are overlooked or multiple instances are merged into a joint segment. UGainS is able to detect and accurately segment many of anomalous objects. However, an overproduction of false positives is observed. In summary, we conclude that there is still plenty of room do develop stronger anomaly instance segmentation and anomalous object detection methods.
V Conclusion
Detecting and accurately segmenting anomaly instances on roads is a significant challenge, requiring an understanding of ’objectness’ without direct training on specific anomaly classes. In this work, we introduced new benchmarks for anomaly instance segmentation and anomalous object detection that integrate three popular anomaly datasets with new instance-wise annotations. This is complemented with an evaluation protocol that treats large and small objects as equally important. Our unified benchmark, termed OoDIS, provides a diverse set of anomalies that vary in size, rarity and the visual conditions in which they are presented. Comprising three datasets, OoDIS constitutes a challenging setting with a substantial number of images and annotation detail. We evaluate the performance of current methods for segmenting anomaly instances and provide an intuition behind the results. Our results show that current techniques struggle particularly with distant and small objects, and with precise segmentation masks. The benchmark results suggest strong opportunities for advancement in the area. As autonomous vehicle technologies continue to evolve, driven by large amounts of data, it remains a challenge to capture all possible real-world situations. Our work addresses the need to evaluate instance segmentation and object detection as a step towards reliable autonomous driving and robot navigation.
Acknowledgment
We acknowledge fruitful discussions with H. Blum, R. Chan, S. Gasperini and S. Rai, as well as a contribution of annotations for Fishyscapes by H. Blum, and help with the benchmark submission from S. Gasperini and S. Rai. M.R. & M.A. acknowledge support by the German Federal Ministry of Education and Research (BMBF) within the junior research group project “UnrEAL” (grant no. 01IS22069). A.N. acknowledges funding by the BMBF project “WestAI” (grant no. 01IS22094D). We thank M. Burdorf, G. Lydakis, C. Schmidt, and others in the lab for discussions and feedback.